# Learning (Approximately) Equivariant Networks via Constrained Optimization

### v1

@@ A.3 Limitations and Further Work
SEGNN - N-body (vanilla) 147,424 1.368$\pm$0.004 0.505$\pm$0.005
SEGNN - N-body (constrained) 937,960 1.512$\pm$0.004 0.527$\pm$0.004
SEGNN - QM9 (vanilla) 1,033,525 118.403$\pm$0.108 11.152$\pm$0.083
SEGNN - QM9 (constrained) 9,869,667 129.039$\pm$0.043 11.402$\pm$0.089
EGNO - MoCap (vanilla) 1,197,176 0.702$\pm$0.007 0.130$\pm$0.000
EGNO - MoCap (constrained) 1,304,762 1.106$\pm$0.004 0.130$\pm$0.000
DGCNN-VN - ModelNet40 (vanilla) 2,899,830 57.826$\pm$0.500 5.954$\pm$0.043
DGCNN-VN - ModelNet40 (constrained) 8,831,612 81.760$\pm$0.169 5.978$\pm$0.006
SchNet - QM9 (vanilla) 127,553 10.389$\pm$0.152 0.944$\pm$0.003
SchNet - QM9 (constrained) 153,543 11.126$\pm$0.155 0.947$\pm$0.002
While our ACE framework is broadly applicable to a range of architectures, it does incur some overhead: during training we introduce extra (non-equivariant) MLP layers, which modestly slow down training (see Table 3). Addressing this, one avenue for future research is to “break” equivariance without adding any auxiliary parameters, e.g., by integrating constraint enforcement directly into existing layers or by designing specialized update rules that can break equivariance while preserving runtime efficiency.
Moreover, our experiments have focused on symmetry groups acting on point clouds, molecules, and motion trajectories. A natural extension is to apply our approach to other symmetry groups and data modalities, such as 2D image convolutions or 2D graph neural networks. Alongside this, co-designing network architectures that anticipate the dynamic tightening and loosening of equivariance constraints could produce models that converge even faster, require fewer samples, or require fewer extra parameters.
A.4 Additional plots for $\gamma$ and $\lambda$
While the main text characterizes only the evolution of the modulation coefficients $\gamma_{i}$, it omits the corresponding dual variables $\lambda_{i}$, which serve as the Lagrange multipliers enforcing each equivariance constraint. Here, we present the full, detailed plots of $\gamma_{i}$ and $\lambda_{i}$ for ModelNet40 (Fig. 8 and Fig. 9) and CMU MoCap (Fig. 10 and Fig. 11), allowing simultaneous inspection of the symmetry relaxation and the accompanying optimization “pressure”. By tracing $\lambda_{i}$ alongside $\gamma_{i}$, one can observe how retained constraint violations accumulate multiplier weight, driving subsequent adjustments to $\gamma_{i}$ and thereby revealing the continuous “negotiation” between inductive bias and training flexibility instantiated by our primal–dual algorithm.
@@ A.4 Additional plots for $\gamma$ and $\lambda$
A.5 Additional Experimental Details
All experiments are implemented in PyTorch [71] using each model’s official code base and default hyperparameters as a starting point. Because several repositories employed learning rates that proved sub-optimal, we performed a grid search for both the primal optimizer and, when applicable, the dual optimizer in our constrained formulation. After selecting the best values, we re-ran every configuration with multiple random seeds. A noteworthy mention is that the test results on the QM9 dataset reported in  Table 2 use different splits for the training, validation and test datasets. For SchNet, we shuffle the dataset obtained from PyTorch Geometric [72] and split the dataset with a 80%/10%/10% train/validation/test ratio. For SEGNN, we use the official splits from [44]. Table 4 lists the learning rates explored and the number of seeds per dataset.
We will make the code and instructions on how to reproduce the experiments publicly available upon acceptance.
@@ A.5 Additional Experimental Details
[…]
All of the experiments were performed on internal clusters that contain a mix of Nvidia RTX A5000, RTX 4090, or A100 GPUs and Intel i9-11900K, AMD EPYC 7742, and AMD EPYC 7302 CPUs. All experiments consumed a total of approximately 1000 GPU hours, with the longest compute being consumed on the QM9 dataset. Runtimes and parameter counts are reported in Table 3.

### v2

@@ A.4 Correcting symmetry mis-specification with ACE
[…]
A.8 Additional plots for $\gamma$ and $\lambda$
While the main text characterizes only the evolution of the modulation coefficients $\gamma_{i}$, it omits the corresponding dual variables $\lambda_{i}$, which serve as the Lagrange multipliers enforcing each equivariance constraint. Here, we present the full, detailed plots of $\gamma_{i}$ and $\lambda_{i}$ for ModelNet40 (Fig.˜9 and Fig.˜10) and CMU MoCap (Fig.˜11 and Fig.˜12), allowing simultaneous inspection of the symmetry relaxation and the accompanying optimization “pressure”. By tracing $\lambda_{i}$ alongside $\gamma_{i}$, one can observe how retained constraint violations accumulate multiplier weight, driving subsequent adjustments to $\gamma_{i}$ and thereby revealing the continuous “negotiation” between inductive bias and training flexibility instantiated by our primal–dual algorithm.
@@ A.8 Additional plots for $\gamma$ and $\lambda$
A.9 Additional Experimental Details
All experiments are implemented in PyTorch [Pas+2019] using each model’s official code base and default hyperparameters as a starting point. Because several repositories employed learning rates that proved sub-optimal, we performed a grid search for both the primal optimizer and, when applicable, the dual optimizer in our constrained formulation. After selecting the best values, we re-ran every configuration with multiple random seeds. A noteworthy mention is that the test results on the QM9 dataset reported in  Table˜2 use different splits for the training, validation and test datasets. For SchNet, we shuffle the dataset obtained from PyTorch Geometric [Fey/Lenssen/2019] and split the dataset with a 80%/10%/10% train/validation/test ratio. For SEGNN, we use the official splits from [brandstetter2022geometric].
In the case of VN-DGCNN, we followed the setup of PennPaper, which trains equivariant models on a sub-sampled resolution of 300 points per point cloud. This differs from the original VN-DGCNN paper [deng2021vn], where 1024 points are used. Moreover, our implementation only tuned the primal and dual learning rates ($\eta_{p},\eta_{d}$), without modifying other components or performing a broader hyperparameter search. Due to these differences, both in input resolution and in the extent of tuning, the reported results are not directly comparable to those in deng2021vn.
Table 5 lists the learning rates explored and the number of seeds per dataset.
The code and instructions on how to reproduce the experiments are publicly available at https://github.com/andreimano/ACE.
@@ A.9 Additional Experimental Details
ModelNet40 { $\{\,0.10,\,0.20\,\}$ { $\{\,1\!\times\!10^{-4},\,1\!\times\!10^{-3}\,\}$ 5
CMU MoCap { $\{\,5\!\times\!10^{-4},\,1\!\times\!10^{-3}\,\}$ { $\{\,5\!\times\!10^{-5},\,2\!\times\!10^{-4},\,1\!\times\!10^{-2}\,\}$ 3
N-Body { $\{\,5\!\times\!10^{-4},\,8\!\times\!10^{-4},\,9\!\times\!10^{-4},\,1\!\times\!10^{-3},\,5\!\times\!10^{-3}\,\}$ { $\{\,5\!\times\!10^{-5},\,1\!\times\!10^{-4},\,8\!\times\!10^{-4}\,\}$ 5
QM9 { $\{\,5\!\times\!10^{-5},\,5\!\times\!10^{-4},\,5\!\times\!10^{-3},\,1\!\times\!10^{-2}\,\}$ { $\{\,5\!\times\!10^{-4},\,1\!\times\!10^{-2},\,1\!\times\!10^{-1}\,\}$ 5
All of the experiments were performed on internal clusters that contain a mix of Nvidia RTX A5000, RTX 4090, or A100 GPUs and Intel i9-11900K, AMD EPYC 7742, and AMD EPYC 7302 CPUs. All experiments consumed a total of approximately 1000 GPU hours, with the longest compute being consumed on the QM9 dataset. Runtimes and parameter counts are reported in Table˜4.